Cant seem to get started on this one. Quadriatic in the unknown.
6/x^2+1/x-2=0 (Substitute u=1/x)

Answer by ewatrrr(24785)


Hi
6/x^2+1/x-2=0 (Substitute u=1/x)
6u^2 + u - 2 = 0
factoring
(3u+2 )(2u-1)=0 Note:SUM of the inner product(4u) and the outer product(-3u) = u
(3u+2 )=0 u = -2/3 1/x = -2/3 x = -3/2
(2u-1)=0 u = 1/2 1/x = 1/2 x = 2
CHECKING our Answer***
6/x^2+1/x-2=0 OR
6 + x - 2x^2 = 0
6 - 3/2 - 18/4 = 0
6 + 2 - 8 = 0